What is the difference between polyethylene glycol and ethylene oxide? How is one toxic and one not? Is it true that amounts of ethylene oxide can be found in polyethylene glycol? And what does it mean when no number follows polyethylene glycol when reading it on an ingredient label ?
1
Expert's answer
2019-04-12T07:49:51-0400

Solution.

Ethylene oxide belongs to the class of dihydric alcohols, and ethylene oxide belongs to the class of epoxides. Ethylene oxide has a mutagenic effect, and ethylene glycol is less toxic. Polyethylene glycol is obtained by subjecting ethylene oxide to alkali catalysts and glycols. Therefore, ethylene oxide can be found in polyethylene glycol. If no number follows polyethylene glycol while reading it on the ingredient label, then the ingredient contains ethylene oxide solution, which gradually turns into polyethylene glycol according to the following reaction equation:

nCH2-O-CH2 + H2O = HO-(CH2-CH2-O)n-H

Answer:

Ethylene oxide belongs to the class of dihydric alcohols, and ethylene oxide belongs to the class of epoxides. Ethylene oxide has a mutagenic effect, and ethylene glycol is less toxic. Polyethylene glycol is obtained by subjecting ethylene oxide to alkali catalysts and glycols. Therefore, ethylene oxide can be found in polyethylene glycol. If no number follows polyethylene glycol while reading it on the ingredient label, then the ingredient contains ethylene oxide solution, which gradually turns into polyethylene glycol according to the following reaction equation:

nCH2-O-CH2 + H2O = HO-(CH2-CH2-O)n-H